Circular No. 9110

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOTICE TO SUBSCRIBERS N. Brickhouse and C. Alcock, Smithsonian Astrophysical Observatory

(SAO), write: “As of 2010 Feb. 1, the SAO will no longer be hosting the Central Bureau for Astronomical Telegrams (CBAT) for the International Astronomical Union (IAU) because SAO has been unable to obtain sufficient funding to sustain the CBAT financially. The CBAT is working, with the support of the IAU, on a transition to a new sponsoring institution. As of this announcement, SAO will not accept new subscription money and expects that the CBAT will announce soon the revised arrangements for subscription payments. During the interim period, SAO will not be able to accept subscriptions.”

EDITORIAL NOTICE The IAU, through its General Secretary, confirms that the CBAT will

continue to operate with the undersigned as Director and that every effort is being made to ensure no disruption of the CBAT’s activities during the transition to its new home. Correspondence now can be sent to the Director via e-mail to dwe green@eps.harvard.edu and via postal mail to: Dr. D. W. E. Green; Room 209; Department of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138; U.S.A.

COMET P/2010 A2 (LINEAR) Z. Sekanina, Jet Propulsion Laboratory, writes that the orientations of

the tail of this comet reported from the observations made between Jan. 7 and 16 (IAUC 9105, 9109; CBET 2134) suggest its formation between January and August 2009. Because of the edge-on projection (with the earth only 2^{o} to 3^{o} below the comet’s orbital plane) it is not possible to decide whether the tail is a product of one or more brief emission events or continuous activity over a period of time. From the tail’s length, the maximum solar radiation pressure acceleration exerted on the dust is estimated at ∼ 0.1 percent of the solar gravitational acceleration, which implies that the smallest dust particles in the tail are ∼ 1 mm in diameter (at an assumed density of 1 g/cm^{3}). From the width of the tail, a lower limit on the normal component of the particle velocity is ∼ 0.1 m/s. An improved estimate can be determined from the tail width around the time of the earth’s crossing the orbital plane on 2010 Feb. 9.

2010 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green